OpenStudy (anonymous):

simplify the expression if not possible write simplified but dont confuse me cuz evry1 has been confusing me on this one =P 2(11d-5)

OpenStudy (anonymous):

Twice 11d is 22d Twice -5 is -10 SO in total 22d - 10.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

2*11d = 22d 2*-5 = -10 22d+(-10) = 22d-10

OpenStudy (anonymous):

\[2(11d-5)=0\]Distribute the 2 \[22d-10=0\]Solve for d (put all other terms on one side of the equal sign) \[22d=10\]\[d=10/22\]Simplify \[d=5/11\]

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

Oh, u wanted d....

OpenStudy (anonymous):

Yea... you need to say "solve for d or x or what ever letter is in the equation."